# Indirect forces in disc-planet interaction

### v1

@@ 1 Introduction
Gravitational coupling between a planet orbiting a central star and a protoplanetary disc in which the planet may be embedded is an important astrophysical problem (GT80). Studying it involves considering a dominant mass, a star, that is orbited by a low-mass object, a planet, with both acting gravitationally on a disc. In this case it is natural to consider disc and planetary dynamics in the frame co-moving with the star since the inner parts of the disc (lying interior to the orbit of the planet) tend to move on Keplerian orbits centered on the star.
@@
There are multiple ways in which indirect forces affect disc-planet interaction. It is well known (GT80) that this gravitational coupling is mediated by the excitation of a density wave that propagates through the disc carrying angular momentum and energy. In the absence of dissipation, the torque exerted by the planet on the disc is entirely carried away by the wave (Papaloizou1984; Gold1989), with the disc ffluid elements experiencing no net loss or gain of the angular momentum. But when dissipation, linear or nonlinear, is present, the wave damps and transfers its angular momentum to the disc, driving its evolution and leading to gap opening. In turn, the torque exerted by the perturbed disc on the planet results in the planetary orbit changing, i.e. orbital migration and eccentricity and inclination evolution. As we will see below, indirect forces enter at some level into the description of all of these phenomena — the effect a planet has on the disc and the back-reaction of the disc on the planetary orbit.
Despite of this, many studies of disc-planet interaction explicitly neglect indirect forces when exploring the perturbation imposed by the planet on a disc or when studying the gravitational back-reaction of the disc on the planetary motion. This approach is usually based on the premise that the indirect forces should not strongly affect a particular aspect of the disc-planet problem and can thus be neglected. However a clear quantitative justification of this expectation is usually not given explicitly, with rare exceptions (Crida2025).
The goal of this work is to provide a quantitative characterization of the impact of indirect forces on different aspects of the disc-planet interaction. Upon presenting our physical setup and numerical methodology in Section 2, we focus on the three key ingredients of the problem. First, we provide a quantitative measurement of the differences in the planet-driven disc perturbation pattern computed with and without the indirect forces (Section 3). Second, we study the impact of indirect forces on the key characteristics of the planet-driven density waves — excitation and deposition torque densities and wave angular momentum flux — and show that neglecting indirect forces can lead to paradoxical results (Section 4). Third, we quantitatively examine the role of indirect forces in modifying the rate of planetary migration [^1] This aspect of disc-planet coupling has been recently explored to some degree by Crida2025; we discuss their work in Section 6. (Section 5).
@@ 2.1 Mathematical foundations
Equations governing the motion of $N>2$ gravitating objects (point masses) in a non-inertial reference frame attached to one of them can be found in many textbooks (e.g. Tremaine2023). They can be easily generalized to a system which includes a continuous massive disc. Focusing on the motion of a disc fluid element at the position vector $\bf R$} in a frame attached to the central star, the governing equation including all indirect terms becomes (e.g. Lin2011)
@@
The first term on the right hand side of (2) is usually interpreted as the acceleration caused by the ‘effective two-body force’ between the star and the planet. However, as shown in Rafikov2025, it is more physically appropriate to consider it as a direct gravitational acceleration due to the star (a part proportional to $M_{\star}$) and the indirect acceleration due to the planet (a part scaling with $M_{\mathrm{p}}$). Indeed, the full indirect acceleration,
@@
This reduction requires quantitative justification, which is not easy to provide in general. For example, if we consider the second and third terms proportional to $M_{\mathrm{p}}$ in the right-hand side of (1), both of them have a general form $GM_{\mathrm{p}}f({\bf R},{\bf R}_{\mathrm{p}})$, where $f$ is some function. Neglecting one of these terms and keeping the other may thus be inconsistent, unless we can explicitly assure that the magnitude of $f$ in one of the terms is small. As a result, in general one should keep (or neglect) both the direct and indirect terms due to a given mass, be it a planet (second and third terms in the right-hand side of (1)) or a disc (fourth and fifth terms proportional to $\Sigma$ in the right-hand side of (1)). Eliminating one of these contributions while keeping the other is not self-consistent, unless one can explicitly demonstrate the smallness of the neglected term. This is in line with the conclusions reached earlier in Crida2022; Crida2025.
In this work we focus on studying a non-self-gravitating disc, which is often a good model of astrophysical discs. This means that when modelling disc dynamics, in equation (1) the direct gravitational term due to other disc elements can be neglected, which is achieved by dropping the fourth term in the right-hand side of the equation (1) responsible for the direct disc gravity. This is formally justified by taking a limit $\Sigma\to 0$. However, in that limit the (fifth) indirect term due to the disc in equation (1) must vanish as well, implying that when the disc self-gravity is neglected, the indirect effect of the disc on its own dynamics must not be considered either (cf. Mittal2015), in agreement with what we said earlier.
This logic can be naturally extended in the opposite direction: when the indirect potential of the disc is dropped from the calculation (which again formally implies a limit $\Sigma\to 0$), ideally one should also not include disc self-gravity (unless one can explicitly demonstrate the smallness of the indirect contribution due to the disc). In other words, the direct gravitational and indirect forces of the disc (as well as any other massive component of the system, e.g. planets) should go hand in hand when exploring its dynamics (Zhu2015): they either need to be considered simultaneously or dropped altogether, which is an important statement of mathematical self-consistency.
@@ 2.2 Physical setup
To model gas dynamics of the disc, we employ an adiabatic, non-barotropic equation of state (EoS) with $\gamma=1.4$, i.e. we allow for radially-varying entropy profile but do not employ additional heating or cooling terms. This thermodynamic setup preserves angular momentum of density waves propagating through the disc in the absence of nonlinear dissipation (see Miranda2020I), simpifying our analysis in Section 4.2. The disc is initially axisymmetric with surface density and temperature (expressed through the sound speed $c_{\mathrm{s}}$) profiles in the power law form:
@@
We consider planetary masses of $M_{\mathrm{p}}=(0.01,0.3)M_{\mathrm{th}}$, where $M_{\mathrm{th}}=M_{\star}h^{3}$ is the characteristic thermal mass delineating the regimes of linear (for $M_{\mathrm{p}}\lesssim M_{\mathrm{th}}$) versus nonlinear (for $M_{\mathrm{p}}\gtrsim M_{\mathrm{th}}$) density waves driven by the planet. As we will see later, nonlinear effects are small initially for $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ but they become pronounced already for $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$, thus our choice of $M_{\mathrm{p}}$ allows us to probe different physical regimes. The planetary potential is softened to avoid singularities on the grid. We use a particular form of the softened potential given by the equation (24) of Cimerman2021, with the softening length set at $r_{s}=0.6H$.
@@ 2.3 Numerical simulations
To quantify the effect of the indirect term in disc-planet interaction we often resort to fully nonlinear hydrodynamic simulations. Most of them are carried out with athena++ (Athenapp2020) in the non-inertial astrocentric (centred on the star), two-dimensional (2D) cylindrical-polar frame. The general setup of our models is similar to that of Cimerman2024b, with the important difference that now we perform runs both including and excluding the indirect term (hereafter IT) in the hydrodynamic equation of motion for the disc fluid, by which we mean the last term in the equation (6).
@@
To provide a useful point of comparison, we also run several simulations of the same system in the barycentric, inertial frame using disco [^2] Specifically, we used the version https://github.com/ajdittmann/Disco. (2016ApJS..226....2D). The domains of these simulations extended from $R=0.1R_{\mathrm{p}}$ to $R=20R_{\mathrm{p}}$, with $N_{r}=2304$ cells and a grid spacing that smoothly transitioned between linear to logarithmic at $R=0.2R_{\mathrm{p}}$. The number of azimuthal zones varied so as to keep $R\Delta\phi\approx\Delta R$ as closely as possible, resulting in $N_{\phi}\approx 2031$. We employed wave damping zones a $R\leq 0.2R_{\mathrm{p}}$ and $R\geq 17R_{\mathrm{p}}$.
@@ 2.4 Linear calculations
To better understand some aspects of disc-planet coupling with or without the indirect term, we also used linear calculations (KP93; RP12; PR12) performed using a linear framework developed earlier in Miranda2019I, Fairbairn2022; Fairbairn2025a. This framework solves numerically a set of linearized (and Fourier decomposed in azimuth) fluid equations for the disc perturbation (see Section 3.1 for some background). The full details of this method can be found in the aforementioned studies.
This framework allows us to provide a detailed analysis of disc perturbation in the limit of very low mass planets, $M_{\mathrm{p}}\ll M_{\mathrm{th}}$, which is performed in Section 3. It also naturally allows us to study important integral characteristics of the planet-driven density waves, such as disc torques and angular momentum flux (see Section 4). This method provides a very powerful analysis tool, cross-checked against direct simulations in a number of disc-planet related problems: impact of non-ideal disc thermodynamics on density wave propagation (Miranda2019I; Miranda2020I; Miranda2020II), spawning of multiple arms by planet-driven density waves (Miranda2019II), negative torque density phenomenon (RP12), torque wiggles (Cimerman2024b), migration and eccentricity evolution of eccentric planets (Fairbairn2025a; Fairbairn2025), etc.
@@ 3 Impact of indirect potential on the surface density perturbation
We start by exploring the effect of the indirect force due to a planet on the structure of the planet-induced perturbation in the disc by comparing surface density perturbation patterns computed with and without the IT and identifying the differences between them. Before moving on to the simulation results, we highlight some important features of indirect forces in the context of disc-planet coupling following from the linear theory (GT80).
For a planet on a circular orbit, one can adopt a Fourier ansatz $\delta x(r,\phi,t)=\delta x_{m}(r)\exp[\mathrm{i}m(\phi-\Omega_{\mathrm{p}}t)]$ for any perturbed disc variable $\delta x$, where $\Omega_{\mathrm{p}}$ is the pattern frequency of the planetary perturbation (equal to its orbital Keplerian frequency) and $m$ is the azimuthal wavenumber. Linearizing the 2D hydrodynamic equations one obtains a very general equation for the evolution of the enthalpy-like variable $\delta h_{m}=\delta P_{m}/\Sigma$ ($P$ and $\Sigma$ being the 2D pressure and surface density) in the form (Miranda2020I)
@@
where $C_{1}$ and $C_{0}$ are functions of the radial distance $R$, with their explicit form being set by the background disk properties (e.g. radial profiles of $\Sigma_{0}$, $c_{\mathrm{s}}$, etc.) and thermodynamic assumptions. The right hand side $\Psi_{m}$ depends on disc properties and thermodynamic assumptions, as well as on the $m$-th Fourier component of the perturbing planetary potential $\Phi_{m}$ and its derivatives. More specific versions [^3] See also Fairbairn2022 for an extension to the case of an eccentric planetary orbit. of the equation (8) have been derived by GT79 for a barotropic (constant entropy) EoS, by Baruteau2008, Tsang2014 for a non-barotropic, adiabatic (radially varying entropy) EoS, by Lee2016 for a locally isothermal EoS (in the homogeneous form), and by Miranda2020I for a disc with $\beta$-cooling. The key point here is that regardless of disc thermodynamics, in linear theory $\Phi_{m}$ affects only the $m$-th component of the disc perturbation.
@@
GT80 showed that a perturbation component $\delta h_{m}$ gets excited by $\Phi_{m}$ at the inner/outer Lindblad resonances where the angular frequency of the disc is $\Omega=\Omega_{\mathrm{p}}m/(m\pm 1)$ (lower/upper sign). In the radial range between the resonances, perturbations are evanescent (exponentially damped) but propagate as a wave outside this region. In a Keplerian disc, the $m=1$ inner Lindblad resonance is absent (it formally resides at $R\to 0$), while the outer Lindblad resonance is located at $R_{1}=2^{2/3}R_{\mathrm{p}}\approx 1.59R_{\mathrm{p}}$. Based on the above description, one expects the $m=1$ component of the perturbation associated with the IT to propagate as a wave outside $R_{1}$ and be exponentially suppressed for $R<R_{1}$.
@@ 3.1 Low-$M_{\mathrm{p}}$ (linear) regime
Looking at Figure 1 one immediately notices that the $\Sigma$ difference has an $m=1$ morphology at every $R\gtrsim R_{\mathrm{p}}$, which is precisely what the linear theory predicts. This difference has a wave-like pattern, which get tightly-wound as $R$ increases, consistent with the linear dispersion relation for density waves in differentially rotating discs (OL02).
@@
A more in-depth analysis of these differences is provided in Figure 2. In panel (a) we show the azimuthal profiles of the surface density perturbation $\delta\Sigma=\Sigma-\langle\Sigma\rangle_{\phi}$ near the density wave maximum at several different radii in the outer disc. We display $\delta\Sigma$ from the runs with the IT (dashed) and without it (solid), normalized by the linear theory prediction for wave amplitude $\delta\Sigma_{\mathrm{lin}}$ (see R02; Cimerman2021). One can see that the density wake profile slowly evolves due to the nonlinear effects present even at such low $M_{\mathrm{p}}$, decaying in amplitude and broadening azimuthally, and the differences between $\delta\Sigma$ computed with and without the IT are rather small in real space but increase with $R$.
Next, in panels (b) and (c) we plot the amplitudes of azimuthal Fourier harmonics of $\delta\Sigma$, $\delta\Sigma_{m}$, for m=16m=1-6 as functions of $R$ inside and outside [^4] Rapid increase of $|\delta\Sigma_{m}|$ near $R_{\mathrm{p}}$ is caused by the quasi-static atmosphere accumulating around the planet and should be ignored. $R_{\mathrm{p}}$, again with and without the IT. In the inner disc, one can see that $|\delta\Sigma_{m}|$ with or without the IT are the same — dashed lines fall on top of the solid ones for all $m\geq 2$. The same is true also for the $m=1$ component, but the more important thing is that $|\delta\Sigma_{1}|$ is very small in amplitude and rapidly decays towards the star. This is very different from $|\delta\Sigma_{m}|$ for $m\geq 2$, which grow as $R$ decreases as a result of approximate conservation of the angular momentum flux (R02). This difference in behaviour is caused by the absence of the $m=1$ inner Lindblad resonance in a Keplerian disc and the lack of direct $m=1$ mode excitation in the inner disc, unlike for the higher $m$ modes. The only reason why $|\delta\Sigma_{1}|$ is non-zero but small for $R<R_{\mathrm{p}}$ is due to the exponentially damped (due to tunneling through the corotation region) $m=1$ perturbation excited at the outer Lindblad resonance at $R_{1}$.

### v3

@@ 1 Introduction
Gravitational coupling between a planet orbiting a central star and a protoplanetary disc in which the planet may be embedded is an important astrophysical problem (Goldreich and Tremaine, 1980). Studying it involves considering a dominant mass, a star, that is orbited by a low-mass object, a planet, with both acting gravitationally on a disc. In this case it is natural to consider disc and planetary dynamics in the frame co-moving with the star since the inner parts of the disc (lying interior to the orbit of the planet) tend to move on Keplerian orbits centered on the star.
@@
There are multiple ways in which indirect forces affect disc-planet interaction. It is well known (Goldreich and Tremaine, 1980) that this gravitational coupling is mediated by the excitation of a density wave that propagates through the disc carrying angular momentum and energy. In the absence of dissipation, the torque exerted by the planet on the disc is entirely carried away by the wave (Papaloizou and Lin, 1984; Goldreich and Nicholson, 1989), with the disc ffluid elements experiencing no net loss or gain of the angular momentum. But when dissipation, linear or nonlinear, is present, the wave damps and transfers its angular momentum to the disc, driving its evolution and leading to gap opening. In turn, the torque exerted by the perturbed disc on the planet results in the planetary orbit changing, i.e. orbital migration and eccentricity and inclination evolution. As we will see below, indirect forces enter at some level into the description of all of these phenomena — the effect a planet has on the disc and the back-reaction of the disc on the planetary orbit.
Despite of this, many studies of disc-planet interaction explicitly neglect indirect forces when exploring the perturbation imposed by the planet on a disc or when studying the gravitational back-reaction of the disc on the planetary motion. This approach is typically based on the premise that the indirect forces should not strongly affect a particular aspect of the disc-planet problem and can thus be neglected. However a clear quantitative justification of this expectation is usually not provided, with rare exceptions (Regály and Vorobyov, 2017; Fairbairn and Rafikov, 2025; Crida et al., 2025).
The goal of this work is to provide a quantitative characterization of the impact of indirect forces on different aspects of the disc-planet interaction. Upon presenting our physical setup and numerical methodology in Section 2, we focus on the three key ingredients of the problem. First, we provide a quantitative measurement of the differences in the planet-driven disc perturbation pattern computed with and without the indirect forces (Section 3). Second, we study the impact of indirect forces on the key characteristics of the planet-driven density waves — excitation and deposition torque densities and wave angular momentum flux — and show that neglecting indirect forces can lead to paradoxical results (Section 4). Third, we quantitatively examine the role of indirect forces in modifying the rate of planetary migration [^1] This aspect of disc-planet coupling has been explored to some degree in the high planetary mass regime by Regály and Vorobyov (2017) and Crida et al. (2025); we discuss these studies in Section 6. (Section 5).
@@ 2.1 Mathematical foundations
Equations governing the motion of $N>2$ gravitating objects (point masses) in a non-inertial reference frame attached to one of them can be found in many textbooks (e.g. Tremaine, 2023). They can be easily generalized to a system which includes a continuous massive disc. Focusing on the motion of a disc fluid element at the position vector $\bf R$} in a frame attached to the central star, the governing equation including all indirect terms becomes (e.g. Lin and Papaloizou, 2011)
@@
The first term on the right hand side of (2) is usually interpreted as the acceleration caused by the ‘effective two-body force’ between the star and the planet. However, as shown in Rafikov (2025), it is more physically appropriate to consider it as a direct gravitational acceleration due to the star (a part proportional to $M_{\star}$) and the indirect acceleration due to the planet (a part scaling with $M_{\mathrm{p}}$). Indeed, the full indirect acceleration,
@@
This reduction requires quantitative justification, which is not easy to provide in general. For example, if we consider the second and third terms proportional to $M_{\mathrm{p}}$ in the right-hand side of (1), both of them have a general form $GM_{\mathrm{p}}f({\bf R},{\bf R}_{\mathrm{p}})$, where $f$ is some function. Neglecting one of these terms and keeping the other may thus be inconsistent, unless we can explicitly assure that the magnitude of $f$ in one of the terms is small. As a result, in general one should keep (or neglect) both the direct and indirect terms due to a given mass, be it a planet (second and third terms in the right-hand side of (1)) or a disc (fourth and fifth terms proportional to $\Sigma$ in the right-hand side of (1)). Eliminating one of these contributions while keeping the other is not self-consistent, unless one can explicitly demonstrate the smallness of the neglected term. This is in line with the conclusions reached earlier in Crida et al. (2022, 2025).
In this work we focus on studying a non-self-gravitating disc, which is often a good model of astrophysical discs. This means that when modelling disc dynamics, in equation (1) the direct gravitational term due to other disc elements can be neglected, which is achieved by dropping the fourth term in the right-hand side of the equation (1) responsible for the direct disc gravity. This is formally justified by taking a limit $\Sigma\to 0$. However, in that limit the (fifth) indirect term due to the disc in equation (1) must vanish as well, implying that when the disc self-gravity is neglected, the indirect effect of the disc on its own dynamics must not be considered either (cf. Mittal and Chiang, 2015), in agreement with what we said earlier.
This logic can be naturally extended in the opposite direction: when the indirect potential of the disc is dropped from the calculation (which again formally implies a limit $\Sigma\to 0$), ideally one should also not include disc self-gravity (unless one can explicitly demonstrate the smallness of the indirect contribution due to the disc). In other words, the direct gravitational and indirect forces of the disc (as well as any other massive component of the system, e.g. planets) should go hand in hand when exploring its dynamics (Zhu et al., 2015): they either need to be considered simultaneously or dropped altogether, which is an important statement of mathematical self-consistency.
@@ 2.2 Physical setup
To model gas dynamics of the disc, we employ an adiabatic, non-barotropic equation of state (EoS) with $\gamma=1.4$, i.e. we allow for radially-varying entropy profile but do not employ additional heating or cooling terms. This thermodynamic setup preserves angular momentum of density waves propagating through the disc in the absence of nonlinear dissipation (see Miranda and Rafikov, 2020b), simpifying our analysis in Section 4.2. The disc is initially axisymmetric with surface density and temperature (expressed through the sound speed $c_{\mathrm{s}}$) profiles in the power law form:
@@
We consider planetary masses of $M_{\mathrm{p}}=(0.01,0.3)M_{\mathrm{th}}$, where $M_{\mathrm{th}}=M_{\star}h^{3}$ is the characteristic thermal mass delineating the regimes of linear (for $M_{\mathrm{p}}\lesssim M_{\mathrm{th}}$) versus nonlinear (for $M_{\mathrm{p}}\gtrsim M_{\mathrm{th}}$) density waves driven by the planet. As we will see later, nonlinear effects are small initially for $M_{\mathrm{p}}=0.01M_{\mathrm{th}}$ but they become pronounced already for $M_{\mathrm{p}}=0.3M_{\mathrm{th}}$, thus our choice of $M_{\mathrm{p}}$ allows us to probe different physical regimes. The planetary potential is softened to avoid singularities on the grid. We use a particular form of the softened potential given by the equation (24) of Cimerman and Rafikov (2021), with the softening length set at $r_{s}=0.6H$.
@@ 2.3 Numerical simulations
To quantify the effect of the indirect term in disc-planet interaction we often resort to fully nonlinear hydrodynamic simulations. Most of them are carried out with athena++ (Stone et al., 2020) in the non-inertial astrocentric (centred on the star), two-dimensional (2D) cylindrical-polar frame. The general setup of our models is similar to that of Cimerman et al. (2024), with the important difference that now we perform runs both including and excluding the indirect term (hereafter IT) in the hydrodynamic equation of motion for the disc fluid, by which we mean the last term in the equation (6).
@@
To provide a useful point of comparison, we also run several simulations of the same system in the barycentric, inertial frame using disco [^2] Specifically, we used the version https://github.com/ajdittmann/Disco. (Duffell, 2016). The domains of these simulations extended from $R=0.1R_{\mathrm{p}}$ to $R=20R_{\mathrm{p}}$, with $N_{r}=2304$ cells and a grid spacing that smoothly transitioned between linear to logarithmic at $R=0.2R_{\mathrm{p}}$. The number of azimuthal zones varied so as to keep $R\Delta\phi\approx\Delta R$ as closely as possible, resulting in $N_{\phi}\approx 2031$. We employed wave damping zones a $R\leq 0.2R_{\mathrm{p}}$ and $R\geq 17R_{\mathrm{p}}$.
@@ 2.4 Linear calculations
To better understand some aspects of disc-planet coupling with or without the indirect term, we also used linear calculations (Korycansky and Pollack, 1993; Rafikov and Petrovich, 2012; Petrovich and Rafikov, 2012) performed using a linear framework developed earlier in Miranda and Rafikov (2019a), Fairbairn and Rafikov (2022, 2025). This framework solves numerically a set of linearized (and Fourier decomposed in azimuth) fluid equations for the disc perturbation (see Section 3.1 for some background). The full details of this method can be found in the aforementioned studies.
This framework allows us to provide a detailed analysis of disc perturbation in the limit of very low mass planets, $M_{\mathrm{p}}\ll M_{\mathrm{th}}$, which is performed in Section 3. It also naturally allows us to study important integral characteristics of the planet-driven density waves, such as disc torques and angular momentum flux (see Section 4). This method provides a very powerful analysis tool, cross-checked against direct simulations in a number of disc-planet related problems: impact of non-ideal disc thermodynamics on density wave propagation (Miranda and Rafikov, 2019a, 2020b, 2020a), spawning of multiple arms by planet-driven density waves (Miranda and Rafikov, 2019b), negative torque density phenomenon (Rafikov and Petrovich, 2012), torque wiggles (Cimerman et al., 2024), migration and eccentricity evolution of eccentric planets (Fairbairn and Rafikov, 2025; Fairbairn and Dittmann, 2025), etc. Our linear calculations use the same physical setup (described in Section 2.2) as our simulations (including the adiabatic EoS, which is implemented via a $\beta$-cooling model with a very large value of $\beta$) allowing a direct comparison between them.
@@ 3 Impact of indirect potential on the surface density perturbation
We start by exploring the effect of the indirect force due to a planet on the structure of the planet-induced perturbation in the disc by comparing surface density perturbation patterns computed with and without the IT and identifying the differences between them. Before moving on to the simulation results, we highlight some important features of indirect forces in the context of disc-planet coupling following from the linear theory (Goldreich and Tremaine, 1980).
For a planet on a circular orbit, one can adopt a Fourier ansatz $\delta x(r,\phi,t)=\delta x_{m}(r)\exp[\mathrm{i}m(\phi-\Omega_{\mathrm{p}}t)]$ for any perturbed disc variable $\delta x$, where $\Omega_{\mathrm{p}}$ is the pattern frequency of the planetary perturbation (equal to its orbital Keplerian frequency) and $m$ is the azimuthal wavenumber. Linearizing the 2D hydrodynamic equations one obtains a very general equation for the evolution of the enthalpy-like variable $\delta h_{m}=\delta P_{m}/\Sigma$ ($P$ and $\Sigma$ being the 2D pressure and surface density) in the form (Miranda and Rafikov, 2020b)
@@
where $C_{1}$ and $C_{0}$ are functions of the radial distance $R$, with their explicit form being set by the background disc properties (e.g. radial profiles of $\Sigma_{0}$, $c_{\mathrm{s}}$, etc.) and thermodynamic assumptions. The right hand side $\Psi_{m}$ depends on disc properties and thermodynamic assumptions, as well as on the $m$-th Fourier component of the perturbing planetary potential $\Phi_{m}$ and its derivatives. More specific versions [^3] See also Fairbairn and Rafikov (2022) for an extension to the case of an eccentric planetary orbit. of the equation (8) have been derived by Goldreich and Tremaine (1979) for a barotropic (constant entropy) EoS, by Baruteau and Masset (2008), Tsang (2014) for a non-barotropic, adiabatic (radially varying entropy) EoS, by Lee (2016) for a locally isothermal EoS (in the homogeneous form), and by Miranda and Rafikov (2020b) for a disc with $\beta$-cooling. The key point here is that regardless of disc thermodynamics, in linear theory $\Phi_{m}$ affects only the $m$-th component of the disc perturbation.
@@
Goldreich and Tremaine (1980) showed that a perturbation component $\delta h_{m}$ gets excited by $\Phi_{m}$ at the inner/outer Lindblad resonances where the angular frequency of the disc is $\Omega=\Omega_{\mathrm{p}}m/(m\pm 1)$ (lower/upper sign). In the radial range between the resonances, perturbations are evanescent (exponentially damped) but propagate as a wave outside this region. In a Keplerian disc, the $m=1$ inner Lindblad resonance is absent (it formally resides at $R\to 0$), while the outer Lindblad resonance is located at $R_{1}=2^{2/3}R_{\mathrm{p}}\approx 1.59R_{\mathrm{p}}$. Based on the above description, one expects the $m=1$ component of the perturbation associated with the IT to propagate as a wave outside $R_{1}$ and be exponentially suppressed for $R<R_{1}$.
@@ 3.1 Low-$M_{\mathrm{p}}$ (linear) regime
Looking at Figure 1 one immediately notices that the $\Sigma$ difference has an $m=1$ morphology at every $R\gtrsim R_{\mathrm{p}}$, which is precisely what the linear theory predicts. This difference has a wave-like pattern, which get tightly-wound as $R$ increases, consistent with the linear dispersion relation for density waves in differentially rotating discs (Ogilvie and Lubow, 2002).
@@
A more in-depth analysis of these differences is provided in Figure 2. In panel (a) we show the azimuthal profiles of the surface density perturbation $\delta\Sigma=\Sigma-\langle\Sigma\rangle_{\phi}$ near the density wave maximum at several different radii in the outer disc. We display $\delta\Sigma$ from the runs with the IT (dashed) and without it (solid), normalized by the linear theory prediction for wave amplitude $\delta\Sigma_{\mathrm{lin}}$ (see Rafikov, 2002a; Cimerman and Rafikov, 2021). One can see that the density wake profile slowly evolves due to the nonlinear effects present even at such low $M_{\mathrm{p}}$, decaying in amplitude and broadening azimuthally, and the differences between $\delta\Sigma$ computed with and without the IT are rather small in real space but increase with $R$.
Next, in panels (b) and (c) we plot the amplitudes of azimuthal Fourier harmonics of $\delta\Sigma$, $\delta\Sigma_{m}$, for m=16m=1-6 as functions of $R$ inside and outside [^4] Rapid increase of $|\delta\Sigma_{m}|$ near $R_{\mathrm{p}}$ is caused by the quasi-static atmosphere accumulating around the planet and should be ignored. $R_{\mathrm{p}}$, again with and without the IT. In the inner disc, one can see that $|\delta\Sigma_{m}|$ with or without the IT are the same — dashed lines fall on top of the solid ones for all $m\geq 2$. The same is true also for the $m=1$ component, but the more important thing is that $|\delta\Sigma_{1}|$ is very small in amplitude and rapidly decays towards the star. This is very different from $|\delta\Sigma_{m}|$ for $m\geq 2$, which grow as $R$ decreases as a result of approximate conservation of the angular momentum flux (Rafikov, 2002a). This difference in behaviour is caused by the absence of the $m=1$ inner Lindblad resonance in a Keplerian disc and the lack of direct $m=1$ mode excitation in the inner disc, unlike for the higher $m$ modes. The only reason why $|\delta\Sigma_{1}|$ is non-zero but small for $R<R_{\mathrm{p}}$ is due to the exponentially damped (due to tunneling through the corotation region) $m=1$ perturbation excited at the outer Lindblad resonance at $R_{1}$.